OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 7 inches and a base that is an equilateral triangle with sides of 6 inches? 63 radical 3 in3 147 in3 126 in3 84 radical 3 in3

OpenStudy (anonymous):

\[V = \frac{s^{2}h\sqrt{3}}{4}\]

OpenStudy (anonymous):

ok, i dont understand like how to plug them in like am lost

OpenStudy (anonymous):

\[V = \frac{(6)^{2}7\sqrt{3}}{4}\]Can you solve it now?

OpenStudy (anonymous):

ok so i know so far that 6^2 is36 but then the 7 radical 3 idk what to do and also the 4

OpenStudy (anonymous):

Alright. \[V = \frac{36 \times 7 \times \sqrt{3}}{4}\]\[V = 9 \times 7 \times \sqrt{3}\]What do you get?

OpenStudy (anonymous):

i got 63 square root 3 which is A

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

ughh thanks am one hard learner when it comes to maths...

OpenStudy (anonymous):

lol. Don't worry. Everyone has trouble with math at one point.

OpenStudy (anonymous):

yea your right :)